题目描述
We have a country containing _N _cities. Each day we choose 2 cities such that there is no road between them and build a road between them. We choose each pair of nonadjacent cities with equal probability. Let X be the number of days before we obtain a connected country. What is the expected value of X? Output the integer part of answer.
Input Format
First line of input as an integer N.
Constraints
$N<=30$
Output Format
Print an integer being the result of the test.
Sample Input 0
3
Sample Output 0
2
Explanation 0
In the first example, first two roads are sufficient for connecting the cities so the answer would be 2.
Sample Input 1
4
Sample Output 1
3
Explanation 1
In the second example if the first three roads of the country are edges of a triangle, then we need a fourth road to make the country connected, otherwise the country would be connected with first three roads. The probability of the former situation is 4/20 (number of triple of roads that make a triangle divided by number of ways we can choose 3 different roads), and the probability of later situation is 16/20. So the result would be 4/204 + 16/203 = 3.2 and since you have to print only the integer part as output, print 3
输入格式
见题面
输出格式
见题面